### v3

# Faster Random Walk-based Capacitance Extraction with Generalized Antithetic Sampling

Abstract
Floating random walk-based capacitance extraction has emerged in recent years as a tried and true approach for extracting parasitic capacitance in very large scale integrated circuits. Being a Monte Carlo method, its performance is dependent on the variance of sampled quantities and variance reduction methods are crucial for the challenges posed by ever denser process technologies and layout-dependent effects. In this work, we present a novel, universal variance reduction method for floating random walk-based capacitance extraction, which is conceptually simple, highly efficient and provably reduces variance in all extractions, especially when layout-dependent effects are present. It is complementary to existing mathematical formulations for variance reduction and its performance gains are experienced on top of theirs. Numerical experiments demonstrate substantial such gains of up to 50% in number of walks necessary as well as in actual extraction times compared to the best previously proposed variance reduction approaches for the floating random-walk.
Index Terms:
VLSI, EDA, Capacitance Extraction, Floating Random Walk, Monte Carlo Estimation, Variance Reduction, Antithetics
I Introduction
Contemporary VLSI designs pose challenges for parasitic capacitance extraction due to the sheer number of conductor metals, and dielectric layers involved. The former can be in the vicinity of hundreds of thousands or even millions, while the latter can be especially numerous and very thin [1]. Advanced technology processes exacerbate design size by imposing additional complexity, for example in the form of non-stratified and conformal dielectrics and other Layout-Dependent Effects (LDE), which themselves heavily influence capacitance extraction [2].
Although classical field solver methods for the capacitance problem do exist, e.g., the Finite Difference (FDM), Finite and Boundary Element Methods (FEM-BEM), the Floating Random Walk (FRW) Monte Carlo solver, offers very attractive advantages comparatively, such as: robustness to geometric complexity thanks to lack of meshing, output locality, i.e., allowing targeted solution on only a subset of domain points, user-tunable accuracy and inherent scalability. Thus, it has been steadily gaining popularity in recent years.
The FRW solver involves an application of Monte Carlo integration on a closed surface around the conductor metal in question [1, 3]. Multiple integrals are estimated concurrently, one for each pairwise capacitance. For consistent Monte Carlo estimation, i.e., given that the Monte Carlo means converge to the true values of the constituent integrals, estimation error is only due to variance. For any finite number of samples this is given by the ratio of the underlying random variable variance over the number of samples. Thus, for a certain target accuracy, in order to expedite extraction, there exist two avenues: either obtain the same number of samples faster, or reduce the inherent variance, so that accuracy is increased with the same number of samples, hence variance reduction.
The specific application of Monte Carlo integration to capacitance extraction amounts to obtaining samples through random walks starting from the integration surface. However, the particular structure of the integral to be estimated poses constraints to exactly what variance reduction methods can be applied, while application itself is also not straightforward. For example, importance sampling can be applied both to choosing points on the integration surface, but also to choosing points on the surface of the first transition domain. The latter can be performed with different importance sampling functions, as showcased by [4] and [3], potentially leading to huge differences in performance benefits. The above discussion shows that variance reduction in the FRW solver requires careful consideration. Ideally, any new variance reduction method should, as much as possible, complement variance reduction mechanisms previously established in the state of the art, so as to reap their benefits already as a baseline.
In this work, we propose such a novel variance reduction approach for FRW-based capacitance extraction, which effectively builds upon previous methods in the literature, but achieves notable further performance gains. Our proposed approach utilizes sampling multiple points on the surface of each first transition domain as starting points of random walks, as does [5], however, crucially, sampled points are guaranteed to have weight values with signs opposite to each other. The method to accomplish this does not strive to fulfill geometric constraints as previous work has attempted, but is inherently data-driven, and relies only on the underlying Green’s function gradient data. Contributions of the paper are as follows:
  • A new algorithm to select starting points for walks on the surface of each first transition domain. The algorithm is impervious to the type of dielectrics (e.g., stratified or not) contained within the domain.
  • Proof that the algorithm reduces variance of the Monte Carlo capacitance estimator, not just in expectation, but for every instantiation of the underlying stochastic weights, under very mild assumptions.
The rest of the paper is organized as follows: In section II, prevalent variance reduction approaches for the FRW are discussed. In section III, a basic overview of current random-walk-based capacitance extraction is given. In section IV, the new algorithm is described and variance analysis is carried out. Numerical results are presented in section V. Finally, conclusions are drawn and relevant proofs are laid out in the appendix.
II Related Work
Variance reduction specific to random walk for capacitance extraction has been an extensively studied theme in the literature in recent years. After preliminary attempts [4], seminal work [1, 3], showed that important performance improvements could be obtained by: i) importance sampling on the surface of the first transition domains, and ii) stratified sampling, regarding the planar faces of the Gaussian surface as strata. These works considered only simple rectilinear conductors. Chains of circuit-connected conductors, ”nets”, complicated things however by requiring net-wide extraction, for which a virtual integration surface was introduced in [6], the Virtual Gaussian Surface (VGS), in place of the Block Gaussian Surface of [3].
More recently, authors in [5] present an approach that enables random walks to share first transition domains, i.e., multiple walks can emanate from a single first transition domain. The very idea of shooting multiple walks from a single transition domain is a central concept in itself and is discussed extensively in [5], reaching the important conclusion that it is multiple walks shot from the first and not so much the subsequent transition domains that contributes substantially to variance reduction. A theoretical variance reduction result is proved, however, the proof requires first transition domains at best with stratified dielectrics and at worst with a single homogeneous dielectric, assumptions, which are not very realistic in practice, especially in the presence of LDE [7, 2]. A specific ”exploration” strategy also has to be employed to determine whether application of the proposed method in any given design will be beneficial or not, since, as shown by their experiments, performance improvements are not always guaranteed with respect to current state-of-the-art.
This approach is the most conceptually similar to ours in the sense that we also propose multiple walks from a single first transition domain. Nonetheless, the way we choose and manipulate starting points of walks on the surface of the first transition domain is quite distinct, both as mathematical formulation and as practical application. The method of [5] chooses points on the first transition domain symmetrically across the boundary of the Gaussian surface, changing the probability density function depending on the number of samples, whereas we choose based on the sign of the weight values corresponding to the points, without requiring geometric symmetry and keeping the probability density function the same for all points of the same transition domain. This is a crucial difference, enabling our approach to offer practical variance reduction guarantees for all designs, regardless of the arrangement of dielectrics contained within the first transition domains, something that eludes [5]. Additionally, because our method is universally better compared to current state-of-the-art, there is no need to decide upon its usage or not in any given extraction.
III Preliminaries
The capacitance extraction problem amounts to calculating all pair-wise capacitances between conductor i, and all other conductors j, all kept at known electric potentials (the mathematical formulation of the problem is equivalent either with conductor i at unit potential and all j, $j\neq{i}$ at zero potentials, or vice versa). With fixed potentials, capacitance is directly related to charge induced on the surface, e.g., of conductor i and according to Gauss’s law this is given by the integral:
[…]
where $G$ is a closed integration surface around conductor $i$, and encompasses only conductor $i$, $\varepsilon(\mathbf{r})$ is the dielectric permittivity function and $\phi(\mathbf{r})$ is the electric potential, both defined at point $\mathbf{r}$ on the integration surface. This is an electrostatic setup governed by a differential equation of elliptic type, for which ample details can be found in the literature [8, 1]. The random walk method employs Monte Carlo integration to compute the above integral, effectively estimating its value via an ensemble average of samples of the integrand at many different points $\mathbf{r}$.
Gaussian surface enclosing conductor iConductor iiConductor jjFirst transition domain
Figure 1: Floating random walk example. Two walks are launched from the integration surface around conductor $i$ and land on the surface of conductor $j$.
The integrand requires the value of the electrostatic potential, which is, however, unknown exactly on the surface $G$. The only sets of points with known potentials are the conductor surfaces. The method takes advantage of a central property of the associated elliptic operator, which is the existence of a propagation (or, equivalently, transition or Poisson) kernel. This is also known as surface Green’s function $P(\cdot)$ in the literature and enables the potential at any point $\mathbf{r}$ to be expressed as a weighted average of the potentials at points $\mathbf{r_{1}}$ on a closed surface enclosing $\mathbf{r}$:
$\phi(\mathbf{r})=\oint_{S_{1}}P(\mathbf{r},\mathbf{r_{1}})\,d\mathbf{r_{1}}$ (2)
For capacitance extraction the most fitting such enclosure is a cube $S_{1}$ centered at $r$, which leads to the notion of a transition cube, also referred to as transition domain [3]. This means that an unknown potential at a point $\mathbf{r}$ of $G$ can be expressed in terms of potentials at further away points $\mathbf{r_{1}}$ lying on the surface of transition domain $S_{1}$ centered at $\mathbf{r}$. And, unknown potentials at $\mathbf{r_{1}}$ can, in turn, be expressed in terms of potentials at even further away points $\mathbf{r_{2}}$ on the surface of transition domain $S_{2}$ centered at $\mathbf{r_{1}}$. Each of these integrals can be estimated in practice by a single sample estimator, i.e., utilizing a single sample on the surface of each respective transition domain according to the surface Green’s function. This procedure of sampling on the surface of each consecutive transition domain can be repeated recursively, for as many steps $n$ as necessary, until the point $\mathbf{r_{n}}$ lies on the surface of a conductor. This recursion, effectively a Markov chain, forms the essence of the FRW method. Substituting (2) into (1), for example once, one can obtain a two-step recursion as follows:
[…]
An illustration of how the method typically works is shown in fig. 1. Again, as in (1), (3) is an integral tackled by Monte Carlo integration, specifically by sampling (Markov) chains of transition domains starting on $G$, creating transition domains $S_{1}$, … , $S_{n}$ and marking what conductor the final transition domains end up at. Note that in (3) the inner integral over the first transition domain $S_{1}$ is quite distinct from integrals over all subsequent transition domains $S_{2}\,...\,S_{n}$, in that it contains the gradient of the surface Green’s function and therefore requires special treatment. Variance reduction methods have been proposed to accelerate handling of the first transition domain and are primarily described in [1, 3]. In summary, importance sampling is applied on both $G$ and $S_{1}$, and the final form of the integral is:
$Q=\oint_{G}\frac{\varepsilon(\mathbf{r})}{F}F\oint_{S_{1}}q(\mathbf{r},\mathbf{r_{1}})w(\mathbf{r},\mathbf{r_{1}})\phi(\mathbf{r_{1}})d\mathbf{r_{1}}d\mathbf{r}$ (4)
where the point $\mathbf{r_{1}}$ is sampled according to $q(\mathbf{r},\mathbf{r_{1}})$ and a weight value $w(\mathbf{r},\mathbf{r_{1}})$ is assigned to it:
[…]
In summary, for any given first transition domain, a point is sampled on its surface according to (5). Furthermore, according to (6), this point corresponds to a weight value with fixed magnitude $\frac{K(\mathbf{r})}{L}$. Only the sign of the weight value can vary, provided the transition domain is kept fixed. The sign of the weight value is determined by the exact position of the point on the surface of the domain according to (6). The random walk method estimates the value of capacitance by averaging many (i.e., hundreds of thousands or even millions) weights of the form (6) corresponding to different transition domains $S_{1}^{(i)}$, all formed with their centers on the integration surface $G$.
IV Generalized Antithetic Sampling
IV-A Motivation and High-level idea
As has been previously investigated in [5], it could be beneficial for variance reduction of the FRW method to sample not just one, but multiple points on the surface of each first transition domain and simultaneously launch random walks from all of them. First, we observe that obtaining multiple samples per first transition domain for variance reduction is very related to the idea of antithetic random variables in the literature of Monte Carlo integration [9, 10]. An antithetic variable (equivalently, sample) strives to be such that the integrand takes on a somehow ”opposite” value compared to that of the original sample, i.e., it takes on a high value when the original sample takes on a low value and vice-versa.
The most counterintuitive aspect of antithetic sampling is that samples involved in the Monte Carlo estimator of an integral can no longer be viewed individually. They can only be viewed in pairs. For example, two antithetic variables prescribe that instead of $2n$ independent samples, $n$ independent pairs of samples are now used for the estimator. The antithetic pairs as units should be independent and identically distributed (iid), whereas the two samples comprising each antithetic pair are not generally independent to each other.
There is no restriction on how to obtain the two samples of each antithetic pair, as long as each of them follows the same marginal distribution. In other words, the joint distribution of the antithetic samples within a pair may be arbitrary, as long as their marginal distribution is the same. Extremely crucial is the fact that, the antithetics Monte Carlo estimator is based on expectation against the common marginal distribution, applied to the pairs as independent units and not against the joint distribution of the pair. The inter-dependency within an antithetic pair is then captured through the correlation of its constituent samples, since only second-moment effects are relevant for variance calculation. The above is an important and largely missed result, generally referred to as generalized antithetic variables [11, 9]. The unbiasedness of the estimator is guaranteed for the sum of the antithetic variables by definition of the method and is completely foreign to such notions as importance sampling weights or sampling from conditional pdfs.
[…]
$\begin{split}\hat{\mu}_{anti}&=\frac{1}{2n}\sum\limits_{i=1}^{n}(f(\mathbf{x})+f(\tilde{\mathbf{x}}))\\ E[\hat{\mu}_{anti}]&=\frac{1}{2n}\sum\limits_{i=1}^{n}2E[f(\mathbf{x})]\\ &=E[f(\mathbf{x})]=\int_{D}p_{marg}(\mathbf{x})f(\mathbf{x})d\mathbf{x}\end{split}$ (7)
Note that the scheme of [5] is conceptually relevant to antithetics, however for selecting the initial (not antithetic) sample, the pdf itself and its support on the surface of the transition domain is changed, since they only initially sample on a fraction of the surface of the domain, not the whole surface. This is the reason why they need to subsequently weigh their samples.
As already mentioned, the samples within each antithetic pair will not be independent, however for second-moment effects such as variance, only their correlation is of interest. In fact, optimal variance reduction is achieved when integrand values are maximally negatively correlated [9, 10]. Our main motivation in this work is that optimal variance reduction through antithetic variables is not explicitly sought after or achieved in [5]. According to theory, maximal negative correlation of the weight values is necessary for optimality, and this amounts to every single weight value obtained from a given first transition domain be paired with a weight value of opposite sign. This is not at all guaranteed with the scheme of [5].
We describe a novel method to obtain sample points on the surface of a given first transition domain, to obtain maximal negative correlation. The first challenge towards this is to address the case of the relevant marginal distribution, i.e., the surface Green’s function not being symmetric, which is often the case in practice, e.g., in the presence of non-stratified dielectrics. The relevant pdf in the integral is $q(\mathbf{r},\mathbf{r_{1}})$, its support set is the whole surface of the cubic transition domain and the weight value $w(\mathbf{r},\mathbf{r_{1}})$ corresponding to each sample is the relevant integrand from (4). To apply generalized antithetic sampling, we need both samples to be marginally distributed according to this exact same pdf. Crucially, we want to deliberately arrange the samples, so that the second sample always corresponds to a weight value with sign opposite to that of the first sample. So, for example, if the first sample corresponds to a negative weight value, we want the second sample to correspond to a positive weight value and vice-versa. By virtue of (6), the magnitude of both of these weight values will be the same for any given first transition domain.
The pair of points on the surface of a given first transition domain with weight values of opposite signs is herafter referred to as an antithetic pair. Then, a way to achieve a single antithetic pair, while keeping the marginal pdf the same, is through sequential independent sampling from this pdf. In other words, the same pdf is utilized repeatedly until the desired weight values of opposite signs are first collected. An outline of the exact algorithm to achieve this is showcased as Algorithm 1.
Algorithm 1 Algorithm to obtain NN generalized antithetic pairs of points on a given first transition domain
INPUT: target number of antithetic pairs N, a transition domain and its associated $q(\mathbf{r},\mathbf{r_{1}}),K(\mathbf{r}),L$
OUTPUT: $2N$ points in space to launch walks from
1:for 1 up to N do
2:  Sample a point $Q$ on the surface of the first transition domain with $q(\mathbf{r},\mathbf{r_{1}})$;
3:  Store the point $Q$ and the corresponding weight value, according to (6), $w$;
4:  repeat
5:   Sample a second point $\tilde{Q}$ on the surface of the first transition domain with $q(\mathbf{r},\mathbf{r_{1}})$;
6:   Store the point $\tilde{Q}$ and the corresponding weight value, according to (6), $\tilde{w}$;
7:  until $w\cdot\tilde{w}<0$
8:  Output the points $Q,\tilde{Q}$} as a single antithetic pair;
9:end for
Algorithm 1 essentially describes a discrete random process. A single trial in the process is sampling a point on the surface of the first transition domain. All trials in the process are obviously independent to each other and the joint and marginal distribution of the sampled points after k trials will be:
$\begin{split}p(\mathbf{x_{1}},...,\mathbf{x_{k}})&=\prod\limits_{i=1}^{k}q(\mathbf{r},\mathbf{r_{i}})\\ p(\mathbf{x_{k}}\rvert\mathbf{x_{1}}...,\mathbf{x_{k-1}})&=p(\mathbf{x_{k}})=q(\mathbf{r},\mathbf{r_{k}})\end{split}$ (8)
The critical idea is that each trial is marginally distributed by construction according to the same pdf, $q(\mathbf{r},\mathbf{r_{1}})$, unique to each first transition domain, as per the prerequisite of generalized antithetic sampling. Note that the outcome of each trial, i.e., whether the weight value is positive or negative, is an additional transformation on the sampled point completely separate to actually sampling on the surface of the domain. Since trials are iid, trials with outcomes that are not of interest to us, (e.g., a string of k2k-2 samples with positive weight after a first sample with positive weight) can simply be discarded and the antithetics method allows this, since it subsequently accounts for the, now dependent, kept trials through covariance calculations on the marginal distribution. The possibly non-trivial joint distribution of the initial and antithetic samples is irrelevant for the antithetics estimator. To our knowledge, this is a novel way of weeding out antithetic samples. If we also incorporate the outcomes of the trials, Algorithm 1 can be described through a Markov chain, specifically shown in the appendix.
[…]
As showcased, the algorithm can be extended in a straightforard manner for the purpose of obtaining not just one, but multiple antithetic pairs per first transition domain. All antithetic pairs produced are also obviously iid as required, since the trials are always iid. The average time for this Markov chain to reach its terminal state is proved in the appendix to be independent from the actual probability of obtaining a positive or negative weight value on the surface of the transition domain. In fact, the algorithm will terminate in 3 steps on average, which makes it quite efficient, considering also that the importance sampling pdf $q(\mathbf{r},\mathbf{r_{1}})$ used for the transitions is only ever computed once for every first transition domain.
Gaussian surface enclosing conductor iConductor iiConductor jjAntithetic sample
Figure 2: Floating random walk with antithetic sampling example. Two walks are initiated from each first transition cube from antithetic samples on their surfaces. The walks starting from antithetic samples are drawn with dashed lines. For the first sample (green lines) both walks land on the surface of conductor $j$, while for the second (purple lines) one finishes on conductor $j$ and the other on conductor ii.
IV-B Variance Computation
In summary, when employing a single antithetic pair, we sample two points on the surface of any given first transition domain according to Algorithm 1 and launch walks from both of these points. Each first transition domain centered on a point of the VGS thus contributes two full random walks and two weight values to the final capacitance calculation. This is similar in concept to [5], with the important difference that the two weight values are guaranteed to be of opposite signs. Although more than one antithetic pairs can be obtained on the surface of each first transition domain, in what follows, for simplicity of exposition, we restrict our analysis to a single antithetic pair per first transition domain.
As also discussed in [5], a potential benefit of starting multiple walks from the same first transition domain, instead of a single walk, is reduction in the number of samples on the VGS necessary for convergence of the FRW method. Taking fewer samples on the VGS means forming fewer first transition domains centered on those samples, and correspondingly fewer matches through a Green’s function precomputation scheme. Specifically, precomputation matching and calculation of $q(\mathbf{r},\mathbf{r_{1}})$ and $K(\mathbf{r})$ only need to take place once for every $N$ random walks.
An even more important potential benefit is variance reduction achieved. In this section, we specifically examine how the mean and variance of the Monte Carlo estimator for capacitance are affected by the correlation among weight values introduced by generalized antithetic sampling (hereafter referred to as GAS). Each transition domain $i$ on the VGS out of $n$ will contribute two full random walks and two weight values to the final capacitance calculation to a total of $2n$.
To form the actual capacitance estimator we need to take into account the potential $\phi_{i}$ of the conductor each walk eventually lands on. The convention we use, typical for capacitance extraction, is set the potential of the master conductor to 0 V, and the potential of all the other conductors to 1 V, the master conductor being where walks start from. With this convention, the estimator essentially comprises weights only from walks that do not land on the master conductor.
$w_{i}$ $=\frac{K_{i}}{2L_{i}}$ $X$ $=\frac{1}{n_{+}}\sum\limits_{i=1}^{n_{+}}w_{i}\cdot\phi_{i}$
$\tilde{w}_{i}$ $=-\frac{K_{i}}{2L_{i}}$ $\tilde{X}$ $=\frac{1}{n_{-}}\sum\limits_{j=1}^{n_{-}}\tilde{w}_{j}\cdot\tilde{\phi}_{j}$
$w_{i}\cdot\tilde{w}_{i}$ $=-w_{i}^{2}\leq 0$ $n_{+}$ $=n_{-}=n,\phi_{i},\tilde{\phi}_{j}\in\{0,1\}$ (9)
The factor of 2 in the denominator of the weights above, serves to make the estimator unbiased in light of the fact that there are now $2n$ random walks or weight values, from each $n$ transition domains on the VGS. If all positive weight values from first transition domains $i$ are arranged in a positive group and all negative values in a corresponding negative group, then $X$ and $\tilde{X}$ are the random variables representing the sample means of these two groups. The equality of group sizes is evident since, for any first transition domain, one positive and one negative weight value will always be produced by construction, see Algorithm 1.
The proposed GAS estimator is simply the sum of the above group sample means:
$\hat{\mu}_{GAS}=F\cdot(X+\tilde{X})=F\cdot\frac{1}{n}(\sum\limits_{i=1}^{n}w_{i}\phi_{i}+\sum\limits_{j=1}^{n}\tilde{w}_{j}\tilde{\phi}_{j})$ (10)
where $F$ is the permittivity integral over the VGS included in (5). GAS is similar to the the Importance Sampling-Stratified Sampling (IS+SS) estimator already used in FRW methodology [1, 3], where $X$ and $\tilde{X}$ can best be thought of as the sample means per positive and negative stratum respectively. The subtle difference is that GAS jointly considers $2n$ weight values for each $n$ first transition domains on the VGS, whereas IS+SS considers only $n$ weight values for the same domains, approximately $n/2$ per stratum. Also, unlike IS+SS, in GAS, $X$ and $\tilde{X}$ are not independent, therefore we need to explicitly examine the mean and variance of the sum of $X$ and $\tilde{X}$ according to the antithetics method [9]:
[…]
The expectation of the sum of group sample means is equal to the sum of the expectations of the sample means, exactly as in the IS+SS case. By contrast, the variance of the sample means is the sum of the sample variances, modified by the covariance of the sample means across the two groups.
Combining with (9), the covariance of $X$ and $\tilde{X}$ can be related to the moments of $w_{i}$, $\tilde{w}_{j}$ as follows:
$\begin{split}Cov[X,\tilde{X}]&=E[X\cdot\tilde{X}]-E[X]\cdot E[\tilde{X}]\\ &=\frac{1}{n^{2}}\sum\limits_{i=1}^{n}E[w_{i}\tilde{w}_{i}]\phi_{i}\tilde{\phi}_{i}\\ &-\frac{1}{n^{2}}\sum\limits_{i=1}^{n}E[w_{i}]E[\tilde{w}_{i}]\phi_{i}\tilde{\phi}_{i}\end{split}$ (12)
The equalities have taken into account that $w_{i}$, $\tilde{w}_{j}$ are uncorrelated for ${i}\neq{j}$, hence $E[w_{i}\tilde{w}_{j}]=E[w_{i}]E[\tilde{w}_{j}]$. This is covariance in expectation. Monte Carlo (or sample) covariance can be estimated from actual weight values $w$, $\tilde{w}$ from (9) and (12) by using sample means in place of actual expectation (as is routinely done for sample variance):
$Cov[X,\tilde{X}]\approx\Delta=\frac{1}{{n^{\prime}}^{2}}\sum\limits_{i=1}^{n^{\prime}}w_{i}\tilde{w}_{i}-\frac{1}{n^{\prime}}(\frac{1}{n^{\prime}}\sum\limits_{i=1}^{n^{\prime}}w_{i})(\frac{1}{n^{\prime}}\sum\limits_{j=1}^{n^{\prime}}\tilde{w}_{j})$ (13)
The above also takes into account that, for non-zero target potentials, only $\phi_{i}\tilde{\phi}_{i}=1$ and $\phi_{i}\tilde{\phi}_{j}=1$ terms survive. Effectively, correlation comes only from antithetic samples where both walks land on conductors other than the master conductor and $n^{\prime}$ is the number of these walks, where $n^{\prime}<n$. We can now prove the following result:
Theorem 1 (Variance Reduction)
Generalized antithetic samples produced by Algorithm 1 lead to maximal negative correlation in (10). Sample variance achieved is always less than the sample variance achieved by the vanilla IS+SS FRW [3] for twice the number of first transition domains, assuming the latter is based on approximately equal total positive vs negative weight values counted from all these domains.
This is a practical result, in the sense that it applies to actual Monte Carlo estimators from finite samples, and not just the theoretical expectations of (10). Furthermore, the dielectrics contained in the first transition domain can be entirely arbitrary. The only remaining assumption to establish universal finite-sample variance reduction is approximate equality of the number of positive and negative weight values over the totality of first transition domains formed on the VGS in the IS+SS case, which is quite reasonable, as discussed in the appendix.
V Numerical Experiments
Algorithm 1 and the associated variance calculation of (13) were implemented within our existing FRW solver and tested in a series of design cases. The test environment used throughout was a Linux machine with Intel Xeon Gold 6326 CPU @ 2.90GHz. All experiments employed a single thread to better capture experienced speedups. Note that the work proposed in this paper does not in any way affect the embarassingly parallel nature of the core FRW method.
The particular variants of FRW estimators compared were:
  • IS+SS: This is used as the baseline estimator, it employs importance sampling and stratified sampling on the first transition domain.
  • SMS-n: This is the multiple-shooting estimator of [5] employing n=2,4,8n={2,4,8} points on each first transition domain selected symmetrically with respect to the VGS.
  • GAS-n: This is the GAS estimator proposed in this paper, employing n=2,4,8n={2,4,8} points on each first transition domain, selected according to Algorithm 1.
As already mentioned, it is entirely possible functionally to take $N$ antithetic pairs per first transition domain with Algorithm 1, instead of just one, and these will be by construction independent to one another. Then, the total sample covariance will be a sum of covariance terms of the form (13) and as such negative as a whole. Furthermore, GAS allows any multiple of 2 as number of samples on the surface of each first transition domain (e.g., 4, 6, 8, 10), as opposed to SMS that allows only powers of 2 (e.g., 4, 8, 16, 32) and thus GAS is more flexible in comparison.
We tested against 9 design cases, roughly divided into two categories: designs without LDE and designs with LDE. The non-LDE design cases (cases 1-4) included simpler structures such as a capacitor and inductor (cases 1-2), as well as more complicated IC designs, e.g., with interposers (cases 3 and 4), all under contemporary techology nodes. The LDE design cases (cases 5-9) were structures of the same vein and technology node as utilized in [2] (5nm) , i.e., metal wires placed above a conductive plane, or between conductive planes, affected specifically by loading, thickness variation and dielectric damage. All relevant effects are applied identically before extraction for all FRW variants. Of these, cases 5-7 did not have a top conductive plane, whereas cases 8-9 did. Aggregate statistics for the test cases are shown in Table I.
TABLE I: Test cases statistics
Case #nets #blocks $blocks_{m}$
1 2 18 18
2 1 3595 3595
3 102 51491 51491
4 20 22711 22711
5 5 116 4
6 5 742 4
7 5 468 4
8 5 571 8
9 5 1209 5
The termination criterion for net extraction in the experiments was always the relative standard error, i.e., the ratio of sample standard deviation over sample mean, to be less than 0.5% (i.e., 0.005). Several independent extractions for each design case were ran to mitigate the inherent stochasticity of the FRW method. The batch size utilized throughout was 1000 to achieve better granularity over possible differences in walks. The LDE cases utilized the machine learning models for Green’s function data from [2] to facilitate extraction.
Besides wall times, another performance metric examined was the number of first transition domains formed until convergence, which we consider to be a more objective metric of variance reduction for these types of approaches, all other things being equal, because it directly measures how many samples are necessary to achieve the same level of accuracy by a Monte Carlo estimator. Apart from the way samples were selected on the surface of the first transition domains, all other infrastructure for our implementation of the variants was identical, including Green’s function data through precomputation and machine learning models. Hence we do not expect hops to differ substantially among the variants and focus on total number of first transition domains. The latter has a straightforward interpretation in terms of walks as well: for IS+SS it is exactly the number of walks, whereas for SMS-n and GAS-n walks it can be found by multiplying the number of first transition domains by $n$.
TABLE II: Results
IS+$S$ SMS-2 GAS-2
Case $C_{tot}(aF/\mu m)$ #first domains time(sec) $C_{tot}(aF/\mu m)$ #first domains time(sec) $C_{tot}(aF/\mu m)$ #first domains time(sec)
1 11.79 604600 18.1 11.47 318900 16.8 11.76 291800 15.9
2 195.9 1798000 89.4 191.59 899000 69.1 196.26 899000 68.1
3 1700317.74 42750100 5046 1651638.25 2645100 672 1696291.44 2642000 600.2
4 10099.06 83884000 10062 9716.43 4849800 2939.9 10042.94 3839200 1874.7
5 25.93 301300 334.2 26.27 166400 270.8 25.9 141100 218.7
6 17.38 159700 413.3 17.25 99100 408.1 17.34 78100 315.3
7 33.54 489300 324.2 32.94 270700 249.8 33.42 245100 221.5
8 15.37 167000 295.1 15.13 116300 290 15.3 82900 207.3
9 24.26 77100 114.8 23.93 54600 108.5 24.29 37900 79.1
TABLE III: Results
SMS-4 GAS-4 SMS-8 GAS-8
Case #first domains time(sec) #first domains time(sec) #first domains time(sec) #first domains time(sec)
1 181600 14.7 145600 13 103700 15.4 73400 11.5
2 450000 48.9 450000 48.8 225000 41.2 225000 40.9
3 2212300 869.2 1659700 635.8 1761100 1225.2 1290800 863.4
4 3035000 3383 2033800 1816.6 1637800 3394.4 1105900 1907.9
5 94700 222.8 71200 158.4 50100 193.7 35900 139.3
6 58400 397.2 39200 257.4 35100 436.7 20000 234
7 151400 203.3 123300 157.4 80300 174.3 61300 128
8 67800 259.8 41800 157.3 40300 268.6 21100 137.1
9 32900 97.8 19000 58.9 20500 102.7 10000 51.5
The main results of experiments for all FRW variants are shown in Tables II and III. The number of first transition domains reported is the sum of first transition domains to extract all nets in the design. Total capacitance $C_{tot}$ reported is the sum of cross capacitances towards conductors of different nets for all nets in the design. Due to space considerations, $C_{tot}$ is only reported for SMS-2 and GAS-2, since extracted values do not substantially differ for the other variants.
For both the multiple-shooting variants (SMS-n and GAS-n) there is a tradeoff between taking more samples on the same first transition domain and forming more (different) first transition domains on the VGS, which also depends on the type of design to be extracted and is also alluded to in [5]. For example, we have experienced substantial performance gains with GAS-2 over IS+SS in virtually every design we have tried, and the same cannot be claimed for GAS-4,8 compared to GAS-2. The best wall times for the largest cases 3 and 4 are not achieved by the GAS-8 variant, but by the GAS-2 and GAS-4 variants respectively. For the SMS schemes, extraction time actually deteriorates when moving from SMS-2 to SMS-4 in cases 3 and 4 and when moving from SMS-4 to SMS-8 in cases 3, 4, 6, 8 and 9. It is hence not straightforward that starting as many walks as possible from each first transition domain is the best approach for every design. We consider additional improvement with respect to GAS-2 and SMS-2 to be highly design-dependent, warranting an ablation study of its own to -ideally- predict whether a given design can benefit from such a choice.
It can be seen that both SMS-n and GAS-n are orders of magnitude faster than IS+SS for the large cases 3 and 4, while also being noticeably faster in the simpler ones (cases 1 and 2). Cases 3 and 4 in particular, quite emphatically affirm the favorable performance results for SMS-n presented in [5] as it achieves large speedups of 7x and 3.5x respectively over IS+SS. Our proposed GAS-2 scheme follows suit with only a very slight edge over SMS-2. For cases 1-3, GAS-2 and SMS-2 are practically equivalent.
On the other hand, there is a distinct advantage of our proposed GAS-n scheme in cases 4-9 over SMS-n, ranging from 9.5% to 50% in terms of first transition domains necessary for convergence, as well as in terms of wall times.
Thorough investigation of why the SMS schemes are lagging in performance is largely outside the scope of this paper. To begin with, SMS variance in [5] uses a different mathematical formulation, where negative correlation among the samples is an emerging, not foundational property. We only describe possible reasons viewed on the basis of covariance (12)-(13), while examining SMS-2 and GAS-2 in particular, to better isolate the essential idea of each method. Under the covariance view, a basic shortcoming of SMS-2 is it has no way of guaranteeing that taking symmetric points with respect to the Gaussian surface will lead to a negative sample covariance in (13). This could happen primarily in cases such as 5-9, where the Green’s function itself exhibits asymmetries, because ambient dielectrics are not necessarily stratified and thus do not profess xy-plane rotational invariance. It could also presumably happen in designs with stratified dielectrics, but now on the top and bottom sides of the Gaussian surface, because there is no symmetry along the z-axis. A second possible and related issue of SMS-2 is that each weight used in the capacitance estimator has an extra multiplicative factor necessary to make the final estimator unbiased. These factors are not fixed across first transition domains and, under the covariance view, could in principle increase final variance. GAS-2 has by construction a fixed weight factor of 0.5, according to (9).
The above reasons are further scrutinized in Table V, where the next to last column reports the percentages of all multiplicative weight factors that differ from 0.5 by more than 0.05. This table also includes the value of $F$ as extra information on the size of the examined designs. A small percentage of symmetric pairs of points on the same first transition domains do in fact have the same sign and produce positive products for SMS-2 as shown in Table V. Note that the sample covariance estimator of (13) is known to exhibit highly non-linear behavior. This means that the first term contributing to $\Delta$ may easily end up being closer to zero (i.e., ”less negative”) than in GAS-2 and hurt overall variance reduction. Furthermore, it can indeed be the case that a significant percentage of weight factors is not close to 0.5 (although not 100%). This is not a full distributional analysis of the weight factors of course, however, it does show that there are both factors close to 0.5 and factors deviating from 0.5, which can, in principle, lead to additional variance of the final SMS-2 estimator. Note that in cases 1-3 where SMS-2 is competitive, either the percentage of positive correlation products is low, or the percentage of non-trivial weight factors is low, or both.
The above comments indicate that the SMS schemes could be affected by factors to which the GAS schemes are agnostic, without compromising efficiency.
TABLE IV: Improvement of GAS-2 over SMS-2
Case $F$
weight pairs
for SMS-2
in domains
1 3.968 0.076 4.95 8.50
2 1.384e+03 0.164 20.89 0
3 5.955e+05 0.237 8.86 0.12
4 7.014e+06 0.43 28.18 20.83
5 1.545 3.54 53.84 15.20
6 1.500 1.63 61.36 21.19
7 2.091 0.2 62.57 9.45
8 0.678 1.24 62.07 28.71
9 0.750 1.12 48.91 30.58
VI Conclusion
We have presented a novel variance reduction approach for Floating Random Walk-based capacitance extraction. The approach builds upon the important idea of shooting multiple walks from each first transition domain, which was recently proposed in the literature and improves upon it by making the selection of points to continue walks from largely data-driven instead of solely geometric. Selection is guided by the sign of the inner product of the gradient of the surface Green’s function and the normal to the Gaussian surface and pairs of points with opposite such signs are obtained by repeatedly sampling from the same pdf through a Markov chain algorithm.
The approach has been shown to reap substantial performance benefits on top of what the currently best variance reduction approaches can afford, especially in scenarios where non-stratified dielectrics make pure geometric point selection less suitable. A thorough investigation of the tradeoff between taking more antithetic pairs on the same first transition domain and forming more first transition domains in the first place remains as future work, in order to possibly decide beforehand what the best number of antithetic pairs per transition domain is for any given design.
Algorithm Analysis
The Markov chain for Algorithm 1 is shown in fig. 3, where $p$ is the probability of getting a positive weight value from a given first transition domain. It is easy to see that it constitutes an absorbing Markov chain with 3 transient states $\mathbf{S,M_{+},M_{-}}$ and a single absorbing state $\mathbf{T}$. The transition matrix is:
$\begin{bmatrix}\mathbf{Q}&\mathbf{R}\\ \mathbf{0}&\mathbf{I}\end{bmatrix}=\begin{bmatrix}0&p&1-p&0\\ 0&p&0&1-p\\ 0&0&1-p&p\\ 0&0&0&1\end{bmatrix}$ (14)
From the theory of absorbing Markov chains [12], it is known that the average number of state transitions to absorption is given by:
$t=(\mathbf{I}-\mathbf{Q})^{-1}\cdot\mathbf{1}=\begin{bmatrix}1&-\frac{p}{1-p}&-\frac{1-p}{p}\\ 0&\frac{1}{1-p}&0\\ 0&0&\frac{1}{p}\\ \end{bmatrix}\cdot\mathbf{1}=3$ (15)
where $\mathbf{1}$ is the vector of all ones. This proves that the Markov chain of Algorithm 1 terminates after 3 transitions on average, irrespective of the actual probability $p$.
SM+M-Tpp1p1-ppp1p1-ppp1p1-p
Figure 3: Markov chain for selecting antithetic points (Algorithm 1). Here $p$ is the probability that the weight value corresponding to a point on the surface of a given first transition domain is positive, $\mathbf{S}$ is the start state, $\mathbf{M_{+},M_{-}}$ are the states after first sampling a point that corresponds to a positive or negative weight value respectively, and $\mathbf{T}$ is the terminal state of the algorithm.
Variance Reduction
From (12), substituting the non-zero target potentials $\phi_{i}\tilde{\phi}_{i}=1$ and $\tilde{w_{i}}=-w_{i}$, covariance can be analyzed as follows:
$Cov_{GAS}=-\frac{1}{n^{2}}(\sum_{k=1}^{n}(E[w_{k}^{2}]-E[w_{k}]^{2}))$ (16)
where for conciseness we have used $n$ in the summation instead of the actual $n^{\prime}$ as the number of non-zero target potentials. Let us examine what happens if we assume an alternative correlation-based estimator, where for a single weight pair the weights do not profess opposite signs. Without loss of generality, we can assume this happens for the first weight pair, i.e., that $\tilde{w_{1}}=w_{1}$. Then:
$Cov_{ALT}$ $=\frac{1}{n^{2}}(E[w_{1}^{2}]-E[w_{1}]^{2})-\frac{1}{n^{2}}(\sum_{k=2}^{n}(E[w_{k}^{2}]-E[w_{k}]^{2}))$
$=\frac{Var[w_{1}]}{n^{2}}-\frac{1}{n^{2}}(\sum_{k=2}^{n}(E[w_{k}^{2}]-E[w_{k}]^{2}))$
$=\frac{2Var[w_{1}]}{n^{2}}-\frac{Var[w_{1}]}{n^{2}}-\frac{1}{n^{2}}(\sum_{k=2}^{n}(E[w_{k}^{2}]-E[w_{k}]^{2}))$
$=\frac{2Var[w_{1}]}{n^{2}}-\frac{1}{n^{2}}(\sum_{k=1}^{n}(E[w_{k}^{2}]-E[w_{k}]^{2}))$
$=\frac{2Var[w_{1}]}{n^{2}}+Cov_{GAS}>Cov_{GAS}$ (17)
since $Var[\cdot]$ is always non-negative. The above can be extended through induction for any number of weight pairs where the opposite sign condition is violated and proves that, in expectation, GAS achieves maximally negative correlation among all variance reduction schemes based on such correlation. Accordingly, it also achieves optimal variance reduction.
The above analysis holds in expectation and a valid question could be: Is there a chance that the actual Monte Carlo covariance from (13) turns out to be positive? Because then the total Monte Carlo variance from (11) would increase and affect algorithm convergence. It turns out that Monte Carlo covariance for GAS-2 is always negative and therefore always reduces Monte Carlo variance:
$\begin{split}\Delta&=\frac{1}{n^{3}}(\sum\limits_{i=1}^{n}nw_{i}\tilde{w}_{i}-\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}w_{i}\tilde{w}_{j})\\ &=\frac{1}{n^{3}}\sum\limits_{i=1}^{n}(nw_{i}\tilde{w}_{i}-\sum\limits_{j=1}^{n}w_{i}\tilde{w}_{j})\\ &=\frac{1}{n^{3}}\sum\limits_{i=1}^{n}[(n-1)w_{i}\tilde{w}_{i}-\sum\limits_{j\neq{i}}^{n}w_{i}\tilde{w}_{j}]\\ &=-\frac{1}{n^{3}}\sum\limits_{i=1}^{n}[(n-1)w_{i}^{2}-\sum\limits_{j\neq{i}}^{n}w_{i}{w}_{j}]\\ &=-\frac{1}{n^{3}}[\sum\limits_{(i,j):i\leq n,j>i}(w_{i}^{2}+w_{j}^{2})-\sum\limits_{(i,j):i\leq n,j>i}2w_{i}{w}_{j}]\\ &=-\frac{1}{n^{3}}\sum\limits_{(i,j):i\leq n,j>i}(w_{i}^{2}+w_{j}^{2}-2w_{i}{w}_{j})\\ &=-\frac{1}{n^{3}}\sum\limits_{(i,j):i\leq n,j>i}(w_{i}-w_{j})^{2}\leq 0\end{split}$ (18)
Note that intermediate results in the proof are easy to derive by expanding and using induction on $n$, e.g., for $n=3$:
$\sum\limits_{(i,j):i\leq 3,j>i}(w_{i}^{2}+w_{j}^{2})=(w_{1}^{2}+w_{2}^{2})+(w_{1}^{2}+w_{3}^{2})+(w_{2}^{2}+w_{3}^{2})$ (19)
In order to compare with vanilla IS+SS fairly, we need to consider double the number of first transition domains for it: in IS+SS, sample variance is calculated every $n$ weight values, obtained from $n$ first transition domains, whereas, in GAS-2, sample variance and covariance are calculated every $2n$ weight values, obtained from $\mathit{n}$ first transition domains. If we further assume that the total number of positive weight values from these first transition domains for IS+SS is approximately equal to the total number of negative weight values (i.e., that the positive and negative strata in IS+SS contain roughly the same number of samples), then (10) can exactly represent IS+SS variance:
$Var[X+\tilde{X}]_{IS+SS}$ $=Var[X]+Var[\tilde{X}$]
$Var[X+\tilde{X}]_{GAS}$ $=Var[X+\tilde{X}]_{IS+SS}+2\Delta,\Delta\leq 0$
Hence, sample variance reduction is always achieved with GAS-2 in the course of an extraction compared to IS+SS. The last assumption of equal number of weights in the positive and negative strata for IS+SS overwhelmingly holds in practice since the first transition domain is generally aligned, by construction, with the vector normal to the VGS and the positive vs negative inner products from (6) average out. This holds even more strongly when the VGS area is larger and as more and more first transition domains are formed on the VGS during extraction.